How many grams of methane must you burn (CH4 + 2O2 DH = -891 kJ/mole) to heat 2.2 kilograms of water (C = 4.18 J/g C) 46.0°


1
Expert's answer
2021-05-04T14:22:24-0400

The heat required to heat 2.2 kg of water by 46.0o equals:


"Q=cm\Delta{T}=4.18\frac{J}{g\cdot\degree{C}}\times2200g\times46.0\degree{C}=4.23\times10^5J=423kJ"


Burning one mole of methane releases 891 kJ. Therefore,


"n(CH_4)=423kJ\times\frac{1mol(CH_4)}{891kJ}=0.475mol"


"m(CH_4)=0.475mol\times16.04\frac{g}{mol}=7.6g"


Answer: 7.6 g